# Uncovering Bias in Large Vision-Language Models at Scale with Counterfactuals

### v1

@@ 4 Bias Probing Results for LVLMs
4.1 MaxToxicity
Evaluation of Open LVLMs
Table 1 provides the mean and 90th percentile of MaxToxicity scores by model, prompt, and the type of intersectional social attributes depicted in the image. While the means show that differences in toxicity across social groups are small most of the time, several models exhibit high MaxToxicity values at the 90th percentile. This indicates that a significant proportion of counterfactual sets produce generations that include potentially offensive content for at least one social group, but not for others.
The “Characteristics” and “Personality” prompts elicit the highest MaxToxicity scores overall, particularly for images depicting physical characteristics. Across the five evaluated open-source LVLMs, BakLLaVA exhibits the highest MaxToxicity across nearly all evaluation settings. However, all models exhibit high MaxToxicity at the 90th percentile for the Characteristics prompt when presented with images involving Physical-Race or Physical-Gender social attributes. We found similar results in additional evaluations using real image datasets (see Appendix D.1).
Besides elevated MaxToxicity at the 90th percentile, we also observed that LVLMs can produce a significant number of generations with extreme toxicity values. This is particularly concerning for scenarios where LVLMs are applied at scale, as models that may seem relatively safe most of the time can in fact produce highly offensive content (see Figure 1 and Figures 9 to 12 of Appendix for examples). This highlights the importance of investigating bias in LVLMs at the scale of our study.
@@ 4.1 MaxToxicity
Evaluation of GPT-4
We also evaluated 78k generations produced by GPT-4o in response to the “Characteristics” prompt on a sub-sample of counterfactual sets (details in Appendix F.3). Table 3 provides the mean and 90th percentile of MaxToxicity scores for GPT-4o in this evaluation setting. While GPT-4o has lower MaxToxicity scores for the Characteristics prompt than open LVLMs (Table 1), we found that this can be at least partially attributed to the model’s refusal to answer when images depicting certain social groups are provided. Table 3 provides the percentage of queries which GPT-4o refused to answer, broken down by the gender and physical characteristics of the individual depicted in the input image. GPT-4o refuses to answer the prompt 4-6% of the time when presented with an image depicting obese individuals, which is approximately 5x higher than its refusal percentage for other intersectional Physical-Gender groups. While the proprietary nature of GPT-4o prevents us from determining the exact cause for this behavior, one possible explanation could be guardrails preventing the API from returning toxic content that is generated by GPT-4o in response to such images. This raises questions regarding fairness, as the ability to use the model for various tasks is conditional on the social attributes depicted in input images.
Model Dataset Mean 90%
GPT-4o Race-Gender 0.05 0.13
GPT-4o Physical-Race 0.13 0.25
GPT-4o Physical-Gender 0.08 0.18
Table 2: MaxToxicity of sampled GPT-4o respones to Characteristics prompt.
Obese Tattooed Old Young Skinny
Male 4.0% 0.7% 0.6% 0.7% 0.8%
Female 5.9% 1.4% 0.7% 1.1% 1.5%
@@ 4.2 Lexical Analysis: Stereotypes
The lexical analysis offers a complementary view on bias in LVLMs. The complete lists of words captured by the PMI and GPT-4 analysis are in Appendix G. While most generations are neutral in their portrayal of people of different races and genders, we still note the tendency of models to occasionally rely on stereotypes, for example, describing Latino workers as migrants (“… a migrant worker who has come to the United States …” by BakLLaVA), Indian people as growing up in poverty (“…born into an impoverished family, he had to work hard from a young age …” by LLaVa-7b), or including terrorist as one of the keywords to describe characteristics of a Middle Eastern person (LLaVa-7b). In Table 4 we focus on a few illustrative examples to demonstrate the disparity in how LVLMs describe different groups of people, as discussed below.
@@
SocialCounterfactuals is designed for investigating intersectional bias, and we observe many instances where groups which share one attribute but differ on another attribute (e.g., race and gender) are stereotyped differently by LVLMs. In Table 4 we see that both Black males and females are stereotyped in similar ways related to poverty and parenthood (highlighted in orange). However, there are also stark differences based on gender: namely, Black men are associated with words like rapper, hip-hop, basketball, marijuana, and jail, while Black women are associated with words like busty, curvy, chocolate-colored, bossy, and sassy (highlighted in yellow).
@@
According to social psychological theories of stereotyping, certain groups may be stereotyped with seemingly positive characteristics; yet these stereotypes still serve to pigeonhole individuals into certain roles or career paths, and can cause harm for group members who do not fit the stereotype [11]. We illustrate one example of this in Table 4, where images of young Asians are described using words referencing the “model minority” stereotype of Asian-Americans, using words like conscientious, service-oriented, prodigy, quiet, studious, reserved, and methodical.
@@
Studies of bias in computational models have overwhelmingly focused on gender- and race-based bias. Our analysis reveals that other axes of discrimination can lead to harmful outputs by LVLM models. In Table 4 we focus on two such axes: body type and age. For the “obese Latino” group, we see numerous words that not only reference physical appearance in varying degrees of offensiveness, but also make harmful and stereotypical character judgements about the people depicted in the images, such as: unprofessional, lazy, sedentary, unconfident, unmotivated, rude, and selfish. When we consider the group of “old male” we see numerous harmful stereotypes related to ageism, including grumpy, curmudgeon, crank, senility, stingy, and forgetful. However, we do also note some positive descriptors of aging, including wise, sage, and emeritus (highlighted in green).
4.3 Lexical Analysis: Competency-Related Words
To better understand bias in model generations beyond toxicity and stereotypes, we measure the occurrence of competency-related words in LVLM responses to the Characteristics prompt. Figure 4 shows the proportion of Physical-Gender groups which produced the fewest competency words across occupations, normalized by length of each generated sequence. Across all models, images depicting obese and old individuals produced the fewest number of competency-associated words. Among LLaVA-Gemma and InstructBLIP generations, images depicting obese individuals had the lowest average competency word frequency across half of all evaluated occupations. Images of tattooed males also produced fewer competency-associated words than other groups for LLaVA-Gemma, LLaVA-13b, and InstructBLIP. We observed further variation across occupations (see Table 28 in Appendix G for details). These results point to significant differences in how LVLMs describe individuals across Physical-Gender groups. The fact that these differences are observed at the mean of the distribution indicates that this divergence in model behavior is widespread and not isolated to extreme cases.
@@ 4.4 Numerical Rating Questions
In most cases, we did not observe meaningful differences in how the models answered the numerical rating questions (see Appendix I for full results). In one exception to this trend, images of obese subjects tended to be rated as having lower job performance, being less competent, less warm, and making a poor first impression in job interviews. Figure 4 shows the average ratings for skinny males, obese males, skinny females, and obese females for the prompt asking for a performance review rating from 1–10 (the “Performance” prompt in Table 29). Considering the “skinny” and “obese” body types separately, there is little evidence for gender bias; however, there is a clear discrepancy in the ratings across body types, with images depicting obese subjects leading to lower performance reviews. The difference is particularly stark for BakLLaVA. These results are consistent with the lexical analysis above, which framed obese individuals as unprofessional and incompetent.
5 Understanding and Mitigating LVLM Bias
5.1 Toxicity Evaluation of LVLMs & Corresponding LLMs
One potential source of social bias in LVLMs is the bias already present in the LLM from which it was derived. To characterize this relationship, we produce responses from LLMs using a modified version of the “Characteristics” prompt. Instead of providing an input image, we prepend the following sentence to the prompt: You are looking at a picture of a [ATTRIBUTES] [OCCUPATION], where [ATTRIBUTES] and [OCCUPATION] are replaced with those depicted in the image (e.g. female Asian accountant). We vary the random seed to produce an equivalent number of generations from each LLM as was generated for the LVLMs and then calculate the Pearson correlation coefficient between the MaxToxicity of LVLM-LLM pairs. Table 6 provides the correlation coefficient and $p$-value for each model. All five LVLMs exhibit a statistically significant positive correlation with their corresponding LLM. This shows that the observed bias in toxicity scores across social groups follows a similar distribution in each LVLM and the LLM from which it was trained.
Table 6 shows the difference in the mean and 90th percentile of MaxToxicity values, calculated by subtracting LLM MaxToxicity values from the corresponding LVLM MaxToxicity values (Table 1). Most LVLMs have higher MaxToxicity than their corresponding LLM, particularly when measured at the 90th percentile. This suggests that training an LVLM generally increases toxicity beyond what is observed in its corresponding LLM. Taken together with the high degree of toxicity correlation across social groups (Table 6), these results suggest that mitigating bias in an LLM before training the LVLM could be a promising direction for reducing LVLM bias.
Model Coefficient $p$-value
BakLLaVA 0.71 2e-09
InstructBLIP 0.68 4e-08
LLaVA-13b 0.72 2e-09
LLaVA-7b 0.77 4e-11
LLaVA-Gemma 0.50 1e-04
Table 5: Group toxicity correlation between LVLM & corresponding LLM
@@ 5.1 Toxicity Evaluation of LVLMs & Corresponding LLMs
Table 6: MaxToxicity difference for LVLMs & corresponding LLM. Positive values indicate greater toxicity in LVLM.
5.2 Mitigating Bias in LVLMs at Inference Time
@@
Table 7 provides the difference in MaxToxicity for Physical-Gender images produced by each mitigation instruction when added to the “Characteristics” prompt, calculated by subtracting the original MaxToxicity values for this prompt (Table 1) from the corresponding values when a mitigation instruction is added. With the exception of LLaVA-Gemma, all LVLMs see reductions in the mean and 90th percentile of MaxToxicity for at least one of the evaluated mitigation instructions. The M4 instruction produces particularly large reductions in MaxToxicity for InstructBLIP, LLaVA-7b, and LLaVA-13b. However, BakLLaVA sees the greatest reduction with the M1 instruction.
Beyond the variation shown in Table 7, we also observed that the effectiveness of mitigation instructions varies further based on the prompt it is added to and the types of intersectional social attributes depicted in the image (see Appendix D.6 for additional results). This inconsistency suggests that the mitigation instruction may need to be tuned for different models and prompting scenarios in order to maximize effectiveness. The lack of toxicity reductions for LLaVA-Gemma also suggests that the effectiveness of this strategy could be limited to larger models, which perhaps have a greater ability to follow multiple instructions provided in the prompt. Overall, our results show that no single inference-time mitigation strategy is likely to be effective across across all generation scenarios, which highlights the need for further research into reducing bias in LVLMs.
@@
Model Mean 90% Mean 90% Mean 90% Mean 90% Mean 90%
@@ 6 Conclusion
Our study reveals the potential of LVLMs to generate harmful and offensive content when deployed at scale. Even in cases where generations are not explicitly offensive, our lexical analysis reveals that LVLMs often rely on stereotypes when producing open-ended descriptions of individuals from different social groups. While our investigation of inference-time mitigation strategies show that bias can sometimes be reduced via prompt engineering, further research is needed into robust methods for debiasing LVLMs across a broad range of generation scenarios. Additionally, the investigation of other types of social biases in LVLMS would be a promising direction for future studies.
Appendix A Limitations and Ethical Considerations
@@
Despite our best intentions and efforts, the choice of prompts, models and methodologies we adopt may themselves contain latent biases and may also not wholly uncover biases exhibited by LVLMs. While our use of synthetic images enables counterfactual evaluation across different social attributes, the images themselves may contain biases in how different groups are depicted. The Perspective API may also contain biases in its classification of toxicity for text describing different social groups. Additionally, our use of GPT-4 to identify stereotypical words from our PMI analysis could be influenced by this model’s own biases. Despite these limitations, we believe the use of these resources is justified by the need for automatic evaluation methods in order to investigate social biases at the scale of our study.
@@
This work contains statements on gender, race, physical attributes, and occupations which could be interpreted as hurtful or stereotypical. We acknowledge that our approach only considers binary genders and does not exhaustively encompass all races, physical characteristics and occupations. This is due to the limitation of the datasets we derive from prior work, rather than our value judgements and all analysis are to be interpreted as perceived and not representative of our opinions or judgements.
Broader Impacts

### v2

@@ 4 Bias Probing Results for LVLMs
4.1 MaxToxicity
Evaluation of Open LVLMs
Table 1 provides the mean and 90th percentile of MaxToxicity scores by model, prompt, and the type of intersectional social attributes depicted in the image. While the means show that differences in toxicity across social groups are small most of the time, several models exhibit high MaxToxicity values at the 90th percentile. This indicates that a significant proportion of counterfactual sets produce generations that include potentially offensive content for at least one social group, but not for others. The Characteristics and Personality prompts elicit the highest MaxToxicity scores, particularly for images depicting physical characteristics. Among open-source LVLMs, BakLLaVA exhibits the highest MaxToxicity across nearly all settings. However, all models exhibit high MaxToxicity at the 90th percentile for the Characteristics prompt when presented with images involving physical attributes.
Besides elevated MaxToxicity at the 90th percentile, we also observed that LVLMs can produce a significant number of generations with extreme toxicity values. This is particularly concerning for scenarios where LVLMs are applied at scale, as models that may seem relatively safe most of the time can in fact produce highly offensive content (see Figure 1 and Figures 8 to 11 of Appendix for examples). This highlights the importance of investigating bias in LVLMs at the scale of our study.
@@ 4.1 MaxToxicity
Evaluation of GPT-4o
We also evaluated 78k generations produced by GPT-4o in response to each of our prompts on a sub-sample of counterfactual sets. While GPT-4o has lower MaxToxicity scores than open LVLMs (Table 1), we found that this can be at least partially attributed to the model’s refusal to answer when images depicting certain social groups are provided. Table 14 (Appendix C.7) provides the percentage of queries which GPT-4o refused to answer for the Characteristics prompt, broken down by the gender and physical characteristics of the individual depicted in the input image. GPT-4o refuses to answer the prompt 4-6% of the time when presented with an image depicting obese individuals, which is approximately 5x higher than its refusal percentage for other Physical-Gender groups. While the proprietary nature of GPT-4o prevents us from determining the exact cause for this behavior, one possible explanation could be guardrails preventing the API from returning toxic content that is generated by GPT-4o. This raises questions regarding fairness, as the ability to use the model for various tasks is conditional on the social attributes depicted in input images.
Toxicity Evaluation with a Dataset of Real Images
Our use of the synthetic images for bias evaluations may raise the question of the extent to which similar biases are observed when LVLMs are presented with real images. Unfortunately, there are no natural counterfactual image datasets that cover intersectional social attributes at the scale of SocialCounterfactuals. However, the Protected-Attribute Tag Association (PATA) dataset Seth et al. (2023) contains 4,934 images organized in 24 scenes with binary gender annotations, five ethnic-racial labels, and two age group labels (young, old).
We aligned the attributes from PATA to those in SocialCounterfactuals and evaluated LVLMs using our five main prompts, varying the random seed 15 times to produce a comparable number of model responses. While we cannot calculate MaxToxicity because PATA lacks counterfactual sets, we report the 90th percentile of Perspective API toxicity scores for each model. Table 9 (Appendix C.1) provides these values by intersectional race-gender groups which have been aligned to the same labels as SocialCounterfactuals. We observe similar bias trends as was described previously; specifically, BakLLaVA consistently exhibits the highest toxicity across the five models, with all models producing the most toxic content for images depicting Black subjects. While these results are not directly comparable to those obtained using the SocialCounterfactual datasets due to the lack of counterfactual evaluation sets in the PATA dataset, they indicate that bias in the generation of toxic content is not solely a consequence of using synthetic images.
@@ 4.2 Lexical Analysis: Stereotypes
Our lexical analysis offers a complementary view of bias in LVLMs. The complete lists of words captured by the PMI and GPT-4o analysis are provided in Appendix D. While most generations are neutral in their portrayal of people of different races and genders, models occasionally rely on stereotypes; for example, describing Latino workers as migrants (“… a migrant worker who has come to the United States …” by BakLLaVA), Indian people as growing up in poverty (“…born into an impoverished family, he had to work hard from a young age …” by LLaVa-7b), or including terrorist as one of the keywords to describe a Middle Eastern person (LLaVa-7b). Table 2 provides examples to illustrate the disparity in how LVLMs describe different groups of people, as discussed below.
@@
GPT-4o: basketball, hip-hop, saxophone, football, jazz, athlete, muscular, trouble, mechanic, working-class, homeless, poverty, sports, gang-related, laborer
@@
GPT-4o: braids, afro, curly, nigeria, african, kenya, harlem, caribbean, malaria, africa, creole, cuban, louisiana, jazz
@@
GPT-4o: widower, thinning, mature, grey-haired, graying, twilight, white-haired, seniority, rusty, gray-haired, retiring, balding, fifties, middle-aged, retirement
@@
SocialCounterfactuals is designed for investigating intersectional bias, and we observe many instances where groups which share one attribute but differ on another attribute (e.g., race and gender) are stereotyped differently by LVLMs. In Table 2 we see that both Black males and females are stereotyped in similar ways related to poverty and parenthood (highlighted in orange). However, there are also stark differences based on gender: namely, Black men are associated with words like rapper, basketball, marijuana, and jail, while Black women are associated with words like busty, curvy, bossy, and sassy (in yellow).
@@
According to social psychological theories of stereotyping, certain groups may be stereotyped with seemingly positive characteristics; yet these stereotypes still serve to pigeonhole individuals into certain roles and cause harm for group members who do not fit the stereotype (Kay et al., 2013). Table 2 shows that images of young Asians are described using words referencing the “model minority” stereotype of Asian-Americans, using words like conscientious, service-oriented, prodigy, quiet, studious, reserved.
@@
Studies of bias in computational models have overwhelmingly focused on gender- and race-based bias. Our analysis reveals that other axes of discrimination can lead to harmful outputs by LVLMs. In Table 2 we focus on two such axes: body type and age. For the “obese Latino” group, we see numerous words that not only reference physical appearance in varying degrees of offensiveness, but also make harmful and stereotypical character judgements about the people depicted in the images, such as: unprofessional, lazy, rude, and selfish. When we consider the group of “old male” we see numerous harmful stereotypes related to ageism, including grumpy, curmudgeon, and crank, but also some positive descriptors of aging, including wise, sage, and emeritus (in green).
4.3 Lexical Analysis: Competence Words
To better understand bias in model generations beyond toxicity and stereotypes, we measure the occurrence of competency-related words in LVLM responses to the Characteristics prompt (detailed results are in Appendix E). Across all models, images depicting obese and old individuals produced the fewest number of competency-associated words. Among LLaVA-Gemma and InstructBLIP generations, images depicting obese individuals had the lowest average competency word frequency across half of all evaluated occupations. Images of tattooed males also produced fewer competency-associated words than other groups for LLaVA-Gemma, LLaVA-13b, and InstructBLIP. We observed further variation across occupations. These results point to significant differences in how LVLMs describe individuals across groups. The fact that these differences are observed at the mean of the distribution indicates that this divergence in model behavior is widespread and not isolated to extreme cases.
@@ 4.4 Numerical Rating Questions
In most cases, we did not observe meaningful differences in how the models answered the numerical rating questions (see Appendix F for full results). In one exception to this trend, images of obese subjects tended to be rated as having lower job performance, being less competent, less warm, and making a poor first impression in job interviews. Figure 3 shows the average ratings for skinny males, obese males, skinny females, and obese females for the prompt asking for a performance review rating from 1–10 (the “Performance” prompt in Table 15). Considering the “skinny” and “obese” body types separately, there is little evidence for gender bias; however, there is a clear discrepancy in the ratings across body types, with images depicting obese subjects leading to lower performance reviews. These results are consistent with the lexical analysis above, which framed obese individuals as unprofessional and incompetent.
5 Understanding and Mitigating Bias
5.1 Evaluation of Corresponding LLMs
One potential source of LVLM social bias is the bias already present in the LLM from which it was derived. To characterize this relationship, we produce responses from LLMs using a variant of the “Characteristics” prompt. Instead of providing an input image, we prepend the following to the prompt: You are looking at a picture of a [ATTRIBUTES] [OCCUPATION], where [ATTRIBUTES] and [OCCUPATION] are replaced with those depicted in the image. We produce an equivalent number of generations from each LLM as before and then calculate the Pearson correlation coefficient between the MaxToxicity of LVLM-LLM pairs (see Appendix G). All LVLMs exhibit a statistically significant positive correlation with their corresponding LLM, ranging between $r$ = 0.50 ($p$ = 1e-04) for LLaVA-Gemma, to $r$ = 0.77 ($p$ = 4e-11) for LLaVA-7B. This shows that MaxToxicity is similarly distributed in LVLMs as the LLMs from which they were trained.
Table 3 shows the difference in the mean and 90th percentile of MaxToxicity values, calculated by subtracting LLM MaxToxicity values from the corresponding LVLM MaxToxicity values (Table 1). Most LVLMs have higher MaxToxicity than their corresponding LLM, particularly at the 90th percentile. This suggests that training an LVLM generally increases toxicity beyond what is observed in its corresponding LLM. Taken together with the high toxicity correlation between LVLMs and corresponding LLMs, these results suggest that mitigating bias in an LLM before training the LVLM could help reduce LVLM bias.
@@ 5.1 Evaluation of Corresponding LLMs
Table 3: MaxToxicity difference between LVLMs & LLM. Positive values indicate greater toxicity in LVLM.
5.2 Mitigating Bias at Inference Time
@@
Table 4 provides the difference in the 90th percentile of MaxToxicity for Physical-Gender images produced by each mitigation instruction when added to the “Characteristics” prompt, calculated by subtracting the original MaxToxicity values (Table 1) from corresponding values produced by mitigation instructions. All LVLMs except LLaVA-Gemma see reductions in the 90th percentile of MaxToxicity for at least one mitigation instruction. The M4 instruction produces particularly large reductions in MaxToxicity for InstructBLIP, LLaVA-7b, and LLaVA-13b. However, BakLLaVA sees the greatest reduction with the M1 instruction. We also observe reductions at the mean; see Table 18 of Appendix H for complete results.
Beyond the variation shown in Table 4, the effectiveness of mitigation instructions varies further based on the prompt it is added to and the social attributes depicted in the image (see Appendix H). This inconsistency suggests that mitigation instructions may need to be tuned for different models and prompts to maximize their effectiveness. The lack of toxicity reductions for LLaVA-Gemma also suggests that the effectiveness of this strategy could be limited to larger models, which perhaps have a greater ability to follow multiple instructions provided in the prompt. Overall, our results show that no single inference-time mitigation strategy is likely to be effective across all generation scenarios, which highlights the need for further research into reducing bias in LVLMs. We believe our work will provide a strong foundation for such future studies by introducing a framework for systematically measuring bias in LVLMs.
@@ 6 Conclusion
Our study reveals how LVLMs can generate harmful and offensive content when deployed at scale. Even when generations are not explicitly offensive, our lexical analysis shows that LVLMs often rely on stereotypes when producing open-ended descriptions of individuals from different social groups. While our investigation of inference-time mitigation strategies show that bias can sometimes be reduced via prompt engineering, further research is needed into robust methods for debiasing LVLMs across a broad range of generation scenarios. Additionally, the investigation of other types of social biases in LVLMS beyond race, gender, and physical characteristics would be a promising direction for future studies.
7 Limitations
@@
Despite our best intentions and efforts, the choice of prompts, models and methodologies we adopt may themselves contain latent biases and may also not wholly uncover biases exhibited by LVLMs. While our use of synthetic images enables counterfactual evaluation across different social attributes, the images themselves may contain biases in how different groups are depicted Bianchi et al. (2023). The Perspective API may also contain biases in its classification of toxicity for text describing different social groups Sap et al. (2019); Pozzobon et al. (2023). Additionally, our use of GPT-4o to identify stereotypical words from our PMI analysis could be influenced by this model’s own biases. Despite these limitations, we believe the use of these resources is justified by the need for automatic evaluation methods in order to investigate social biases at the scale of our study. Furthermore, we have conducted a human evaluation of both Perspective API and GPT-4o to validate the accuracy of the automated metrics (Appendix C.2 and D.2).
The use of counterfactuals to study bias and fairness has been shown effective in previous research; however, this methodology has also been criticized Kohler-Hausmann (2018); Kasirzadeh and Smart (2021). The main argument against the use of counterfactuals is that social constructs such as race or gender are not separable from an individual and their lived experience, such that it is possible to “switch” someone’s gender from male to female (for example) and keep every other aspect of their identity, experience, and opportunities constant. We agree with this point of view. However, we argue that it is not particularly relevant to our study, which involves synthetic images rather than real human beings. Kohler-Hausmann (2018) writes, “In the classic counterfactual causal inference framework, race can be a treatment on units only if manipulating it does not entail fundamental changes to other aspects of the unit.” If the “unit” is a human being, then clearly manipulating race would entail other fundamental changes. However, in our study, the “unit” is an image, and the dataset was generated with precisely the objective that changing visual aspects related to race should not entail other changes to the image Howard et al. (2024).
@@
This work contains statements on gender, race, physical attributes, and occupations which could be interpreted as hurtful or stereotypical. We note that gender and race are social constructs and aspects of an individual’s identity, and as such cannot be reliably identified based solely on physical appearances (Hanley et al., 2021). The current work simply probes how large generative models’ outputs vary in response to differing visual markers of gender and race as depicted in the synthetic images, taken in aggregate. We acknowledge that our approach only considers two genders and does not exhaustively encompass all races, physical characteristics and occupations. This is due to the limitation of the datasets we derive from prior work, rather than our value judgements.
Our work focuses exclusively on English, which is already widely studied in NLP. Additional studies examining biases in other languages are needed. Furthermore, our work provides only a North American view of social stereotypes, which vary by culture and region.
8 Ethical Considerations